3). I need to make a 0.50M Tris buffer with a pH of 8.75. I have access to Tris (a weak acid) and the sodium salt of Tris (think A-). The Ka of Tris is 8.5x10-9 and the pKa of Tris is 8.07. What are the [HA] and [A-] concentrations of this 0.50M buffer at pH 8.75?  

Part B: [A-] = _____
Answer #1

[HA] + [A-] = 0.50M

pH = pKa + log([A-]/[HA])

8.75 = 8.07 + log([A-]/[HA])

log([A-]/[HA]) = 0.68

[A-]/[HA] = 100.68 = 4.7863

[A-] = 4.7863 × [HA]

[HA] = 0.50M/(4.7863+1) = 0.0864M

[HA] = 0.0864M (Answer)

[A-] = 0.0864 × 4.7863 = 0.4136 M (Answer)